OpenStudy (anonymous):

5x+7y=-2, 2x-7y=9 solve by using an algebraic method.

OpenStudy (anonymous):

5x+7y=-2, 2x-7y=9 solve by using an algebraic method.

OpenStudy (anonymous):

algebraic method means ANY method..

OpenStudy (anonymous):

Solve using an algebraic method: 8x-5y=-17 and -2x+y=6